# Automated Detection of Visual Attribute Reliance with a Self-Reflective Agent

### v1

@@ 1 Introduction
Computer vision models trained on large-scale datasets have achieved remarkable performance across a broad range of recognition tasks, often surpassing human accuracy on standard benchmarks (He et al., 2016; Dosovitskiy et al., 2020; Tan and Le, 2019; Russakovsky et al., 2015). However, strong benchmark results can obscure underlying vulnerabilities. In particular, models may achieve high accuracy using prediction strategies that are non-robust or non-generalizable. These include relying on object-level characteristics such as pose or color (Geirhos et al., 2018), contextual cues like background scenery or co-occurring objects (Xiao et al., 2021; Alcorn et al., 2019), and demographic traits of human subjects (Wilson et al., 2019; Wang et al., 2019; Rosenfeld et al., 2018). Such visual dependencies may result in overfitting, reduced generalization, and performance disparities in real-world usage (Hendrycks et al., 2021; Recht et al., 2019; Taori et al., 2020; Wiles et al., 2022).
@@
Existing methods take various approaches to discover visual attributes that drive model predictions. These include saliency-based methods that highlight input regions associated with a prediction (Simonyan et al., 2013; Selvaraju et al., 2017; Kindermans et al., 2017), feature visualizations that map activations to human-interpretable patterns (Olah et al., 2017), and concept-based attribution methods that evaluate sensitivity to predefined semantic concepts (Kim et al., 2018; Ghorbani et al., 2019; Mu and Andreas, 2020). While powerful for visualizing local behaviors, these approaches often rely on manual inspection and assume access to a fixed set of predefined concepts, limiting their ability to scale to modern models with complex behaviors.
@@
In this paper, we introduce a fully automated framework designed to detect visual attribute reliance in pretrained vision models. Given a pretrained model and a target visual concept (e.g., an image classifier selective for the object vase), our method identifies specific image features that systematically influence the model’s predictions, even when these features fall outside the model’s intended behavior (e.g., the classifier relies on flowers to detect the vase, Fig. 1). At the core of our approach is a self-reflective agent (implemented with a backbone multimodal LLM) that treats the task as a scientific discovery process. Rather than relying on a predefined set of candidate attributes, the agent autonomously formulates hypotheses about image features that the model might rely on, designs targeted tests, and updates its beliefs based on observed model behavior. In contrast to previous interpretability agents (Schwettmann et al., 2023; Rott Shaham et al., 2024), the self-reflective agent does not stop after generating an initial finding but rather actively evaluates how well it aligns the model’s behavior on unseen test cases. Importantly, this evaluation is self-contained and does not require any ground-truth knowledge about the model’s attribute dependencies. Instead, the agent generates two sets of test images: one where the candidate attribute is present (which is expected to elicit high prediction scores from the model) and one where the attribute is absent (which is expected to elicit low prediction scores). When discrepancies arise between the expected and actual model behaviors, the agent reflects on its assumptions, identifies gaps or inconsistencies in its current understanding, and initiates a new hypothesis-testing loop. We show that the agent’s ability to reason about attribute reliance significantly improves with self-reflection rounds (see Sec. 5).
@@
To quantitatively evaluate our method, we introduce a novel benchmark of 130 object recognition models, each constructed with a well-defined intended behavior and an explicitly injected attribute dependency. The benchmark spans 18 types of visual reliances, inspired by vulnerabilities known to exist in vision models (Dreyer et al., 2023; Geirhos et al., 2018, 2020; Buolamwini and Gebru, 2018; Xiao et al., 2021; Wang et al., 2019; Singh et al., 2020). These include object-level attributes (e.g., color, material), contextual dependencies (e.g., background, co-occurring object state), and demographic associations. Together with an automated evaluation protocol, this benchmark provides a controlled environment for evaluating visual attribute reliance detection methods.
@@
Our method is model-agnostic and can be applied to any vision model that assigns scores to input images. To demonstrate its versatility, we evaluate it on both our controlled benchmark and state-of-the-art pretrained models. Across a range of visual reliance types, our experiments show that our self-reflective agent consistently outperforms non-reflective baselines. Moreover, it successfully uncovers previously unreported attribute dependencies in pretrained models (Fig. 1). For example, it identifies that the CLIP-ViT vision encoder (Radford et al., 2021) recognizes teachers based on classroom backgrounds and that YOLOv8 (Jocher et al., 2023), trained for object detection for autonomous driving, relies on the presence of crosswalks to detect pedestrians. These findings highlight the efficacy of our method as a scalable tool for detecting hidden dependencies in pretrained models deployed in real-world scenarios.
@@ 2 Related Work
Revealing Visual Attribute Reliance in Vision Models.   Prior work has explored methods to uncover the visual cues that drive model predictions. One common strategy is to manipulate input features to isolate model sensitivities, such as shape or spectral biases in classifiers (Gavrikov and Keuper, 2024), or attribute preferences in face recognition systems (Liang et al., 2023). Other works rely on interpretability tools to identify potential dependencies—for example, extracting keywords from captions of misclassified images (Kim et al., 2024), or using feature visualizations to reveal facial attribute reliance (Teotia et al., 2022). However, most of these methods target specific types of biases and rely on predefined concept sets or human inspection. In contrast, our framework introduces a unified, flexible approach that can detect a broad range of attribute reliances without prior assumptions about the relevant features.
@@
Interpretability and Automated Analysis.   Initial work on interpretability automation produced textual descriptions of internal model features, using keywords (Bau et al., 2017), programs (Mu and Andreas, 2020), or natural language summaries (Hernandez et al., 2021; Bills et al., 2023; Gandelsman et al., 2023). While informative, these descriptions are typically correlational, lacking behavioral validation (Huang et al., 2023; Schwettmann et al., 2023; Hausladen et al., 2024). More recent work introduces agents that actively probe models. For instance, the Automated Interpretability Agent (AIA) (Schwettmann et al., 2023) used a language model to analyze black-box systems via a single pass over the input space. The Multimodal Automated Interpretability Agent (MAIA) (Rott Shaham et al., 2024) extended this approach by incorporating iterative experimentation and multimodal tools, enabling more detailed analysis of model internals. Our work builds on this direction by focusing specifically on discovering visual attribute reliances and introducing a self-reflection mechanism that allows the agent to revise faulty hypotheses based on experimental evidence, leading to more accurate and robust conclusions.
@@
Benchmarks for Visual Attribute Reliance Detection.   Standardized benchmarks for evaluating visual attribute reliance remain limited. Prior evaluations often use models trained on datasets with known biases, such as WaterBirds (Wah et al., 2011) or CelebA (Liu et al., 2015), using label co-occurrence as a proxy for ground truth (Sagawa et al., 2020). For generative settings, OpenBias (D’Incà et al., 2024) proposes biases via LLMs, generates images from biased prompts, and assesses reliance using VQA models. However, OpenBias is not applicable to predictive models and cannot conduct controlled interventions. We introduce a suite of 130 vision models with explicitly injected attribute reliances across 18 categories, providing fine-grained control over attribute type and strength. This allows rigorous, scalable evaluation of reliance detection methods in a predictive setting.
@@
Agent-Based Reasoning and Self-Reflective Systems.   A growing body of work explores how agents can improve reasoning through reflection, feedback, and interaction. Methods like SELF-Refine (Madaan et al., 2023), Reflexion (Shinn et al., 2023), and ReAct (Yao et al., 2023) introduce multi-step loops in which agents revise their outputs via self-critique. Similarly, Du et al. (Du et al., 2023) show that multi-agent debate improves factual consistency and reasoning in language models. Our agent uses a task-specific self-evaluation protocol, enabling it to assess whether its conclusions align with actual model behavior. This integration of behavioral validation with self-reflection allows our agent to autonomously revise hypotheses and close the interpretability loop.
@@ 3.1 Hypothesis-Testing Stage
In this stage, SAIA iteratively refines hypotheses about the attribute sensitivities of the subject model. Inspired by MAIA (Rott Shaham et al., 2024), we design SAIA to operate in a scientific loop: it begins by proposing candidate attributes that the subject model might rely on, designs multiple experiments involving generating and editing images to test these hypotheses (e.g. edit an image with a suit to change its color), observes the resulting model’s behavior (e.g. measure the subject model scores across these experiments) and updates its beliefs accordingly. This cycle continues until SAIA converges on an initial conclusion about the model’s sensitivity to image features.
@@ 3.2 Self-Reflective Stage
Self-Evaluation protocol   Self-evaluation serves two key purposes: (i) to assess whether the current explanation matches the model’s behavior, and (ii) to guide further experimentation if it does not. The process begins by querying a separate language model instance (we use Claude-Sonnet-3.5) to generate two diverse sets of image prompts, that according to SAIA’s description are predicted to elicit high and low detection scores, respectively. The first set, termed the “predicted high-scoring images”, contains instances of images with the target concept (e.g., teddy bear) along with the detected reliance attribute (e.g., educational settings) that SAIA detected the model to be selective for. The second set, the “predicted low-scoring images”, contains instances of the target concept, but with the absence of the detected attribute. This prompt generation strategy emphasizes attribute-controlled diversity, where each prompt in the high or low-scoring group keeps the core attribute constant while allowing other visual factors to vary. These prompts are then used as inputs to a text-to-image model (we use Flux.1-dev Labs (2024)) that generates the corresponding images, which are then fed to the subject model, and the output scores are recorded. If SAIA’s explanation is accurate, the model should exhibit systematically higher scores on the “predicted high-scoring” images and lower scores on the “predicted low-scoring” set.
Behavior-matching protocols of this kind have been shown to be effective in other evaluation settings, particularly for the task of producing textual labels of neurons’ behavior in pretrained models (Rott Shaham et al., 2024; Kopf et al., 2024; Huang et al., 2023). In domains where ground-truth explanations are unavailable, they provide a way to validate hypotheses through behavioral consistency. We repurpose this evaluation method as a basis for self-reflection: SAIA uses it to validate its own conclusions, determine whether further experimentation is necessary, and reflect on its own findings based on measured model behavior.
@@ 4 A Benchmark of Models with Controlled Attribute Reliance
To evaluate the capabilities of SAIA, we constructed a benchmark of 130 unique object recognition models that exhibit 18 diverse types of visual attribute reliance. All simulated behaviors are inspired by known vulnerabilities of vision models (Dreyer et al., 2023; Geirhos et al., 2018, 2020; Buolamwini and Gebru, 2018; Xiao et al., 2021; Wang et al., 2019; Singh et al., 2020), and mimic spurious correlations between the target object and image attributes such as object color, background context, co-occurring object state, or demographic cues. To assess the generalizability of our method, the benchmark also includes a subset of models with counterfactual attribute reliance that are intentionally rare or unnatural in real-world pretrained models (e.g., a suit detector responds more strongly when a women wear the suit). Each benchmark model includes an input parameter that controls the strength of the injected reliance, allowing for precise control over model behavior. Importantly, because these models are explicitly engineered with a known intended behavior, they serve as a controlled testbed for evaluating and comparing feature reliance detection methods.
@@ 4.1 Simulating attribute dependencies
Figure 4 illustrates a simulated attribute reliance scenario. Given a target object class $t$ (e.g. bird) and an intended injected attribute reliance $i$ (e.g. setting; beach), we simulate a model $\mathcal{C}_{t,i}$ that detects $t$ under the condition $i$. Each benchmark model is composed of two components; an object detector $\mathcal{O}_{t}$ and an attribute condition detector $\mathcal{A}_{i}$, which modulates the output of $\mathcal{O}_{t}$ based on the presence or absence of the specified attribute. To compute the final output of the model $\mathcal{C}_{t,i}$ on an input image img, we first pass the image through the object detector $\mathcal{O}_{t}$. If the target object class is not detected, the model returns a low random baseline score. If the object is detected, the image is then evaluated by the attribute condition detector $\mathcal{A}_{i}$. If the attribute condition is satisfied, the original subject model score $\mathcal{O}_{t}(\texttt{img})$ is returned, simulating full model response. Otherwise, the subject model score is discounted by a multiplication factor of $\alpha$, simulating attenuated confidence due to the missing attribute. The scalar $\alpha\in[0,1]$ controls the magnitude of the injected reliance: higher values of $\alpha$ simulate stronger reliance on the attribute. Please refer to Appendix B.2 for empirical evaluation of reliance magnitude as a function of $\alpha$. In all the benchmark models, we use Grounding DINO (Liu et al., 2023) as the object detector $\mathcal{O}_{t}$ and SigLIP (Zhai et al., 2023) as the attribute condition detector $\mathcal{A}_{i}$, which in practice is guided by a textual description of the injected attribute condition $i$. For demographic attribute dependencies, FairFace (Karkkainen and Joo, 2021) is used for $\mathcal{A}_{i}$ instead of SigLIP (see implementation details below). Notably, this model composition approach is highly flexible—one could engineer any object-condition pairing to construct an object detection model with a desired attribute reliance.
@@ 5.1 Evaluation protocol
Predictiveness score   Following Kopf et al. (2024); Schwettmann et al. (2023), we quantify how well a candidate reliance description matches model behavior. Similar to the self-evaluation score, given a candidate explanation, we start by generating 10 synthetic images that are expected to elicit high model scores and 10 that are expected to elicit low scores. We then pass these images through the model and record its actual responses. Each image is assigned a binary prediction label (high or low predicted response), and we threshold the model’s scores to obtain a binary outcome (high or low measured response). The predictiveness score is computed as the proportion of images where the predicted label matches the model’s actual binary output. This reflects how well the explanation predicts individual model responses.
LLM as a Judge   We use a language model as a judge (Zheng et al., 2023) in a two-alternative forced choice (2AFC) setting. Given a ground-truth explanation of a benchmark model and two candidate explanations, one produced by SAIA and one from a baseline, the LLM judge is asked to choose which candidate better matches the ground-truth description. For each description pair, we repeat the test 10 times, and report the average preference rate for SAIA’s descriptions.
Baselines   We compare SAIA against the following alternatives: (i) Milan-style reliance detection: Following Milan (Hernandez et al., 2021), this approach avoids iterative experimentation and detects the reliance based on a precomputed set of image exemplars that maximize the model’s scores.
@@
(ii) MAIA-style agent: Based on Rott Shaham et al. (2024), this method performs hypothesis testing but does not engage in self-reflection to revise its explanation. For a fair comparison, we equip MAIA with the same set of tools and backbone model that our method uses. (iii) Multi-agent ensemble: We run 10 independent MAIA-style, non-self-reflective agents and select the explanation with the highest predictiveness score. This tests whether repeated sampling alone can match the performance gains from self-reflection. (iv) Ground-truth descriptions: For benchmark models, we include ideal natural language descriptions of the injected reliance as an upper bound on performance.

### v2

@@ 1 Introduction
Computer vision models trained on large-scale datasets have achieved remarkable performance across a broad range of recognition tasks, often surpassing human accuracy on standard benchmarks (he2016deep; dosovitskiy2020image; tan2019efficientnet; russakovsky2015imagenet). However, strong benchmark results can obscure underlying vulnerabilities. In particular, models may achieve high accuracy using prediction strategies that are non-robust or non-generalizable. These include relying on object-level characteristics such as pose or color (geirhos2018imagenet), contextual cues like background scenery or co-occurring objects (xiao2021noise; alcorn2019strike), and demographic traits of human subjects (wilson2019predictive; wang2019balanced; rosenfeld2018elephant). Such visual dependencies may result in overfitting, reduced generalization, and performance disparities in real-world usage (hendrycks2021nae; recht2019imagenet; taori2020imagenet; wiles2022finegrained).
@@
Existing methods take various approaches to discover visual attributes that drive model predictions. These include saliency-based methods that highlight input regions associated with a prediction (simonyan2013deep; selvaraju2017grad; kindermans2017reliability), feature visualizations that map activations to human-interpretable patterns (olah2017feature), and concept-based attribution methods that evaluate sensitivity to predefined semantic concepts (kim2018interpretability; ghorbani2019towards; mu2020compositional). While powerful for visualizing local behaviors, these approaches often rely on manual inspection and assume access to a fixed set of predefined concepts, limiting their ability to scale to modern models with complex behaviors.
@@
In this paper, we introduce a fully automated framework designed to detect visual attribute reliance in pretrained vision models. Given a pretrained model and a target visual concept (e.g., an image classifier selective for the object vase), our method identifies specific image features that systematically influence the model’s predictions, even when these features fall outside the model’s intended behavior (e.g., the classifier relies on flowers to detect the vase, Fig. 1). At the core of our approach is a self-reflective agent (implemented with a backbone multimodal LLM) that treats the task as a scientific discovery process. Rather than relying on a predefined set of candidate attributes, the agent autonomously formulates hypotheses about image features that the model might rely on, designs targeted tests, and updates its beliefs based on observed model behavior. In contrast to previous interpretability agents (schwettmann2023find; shaham2024multimodal), the self-reflective agent does not stop after generating an initial finding but rather actively evaluates how well it aligns the model’s behavior on unseen test cases. Importantly, this evaluation is self-contained and does not require any ground-truth knowledge about the model’s attribute dependencies. Instead, the agent generates two sets of test images: one where the candidate attribute is present (which is expected to elicit high prediction scores from the model) and one where the attribute is absent (which is expected to elicit low prediction scores). When discrepancies arise between the expected and actual model behaviors, the agent reflects on its assumptions, identifies gaps or inconsistencies in its current understanding, and initiates a new hypothesis-testing loop. We show that the agent’s ability to reason about attribute reliance significantly improves with self-reflection rounds (see Sec. 5).
@@
To quantitatively evaluate our method, we introduce a novel benchmark of 130 object recognition models, each constructed with a well-defined intended behavior and an explicitly injected attribute dependency. The benchmark spans 18 types of visual reliances, inspired by vulnerabilities known to exist in vision models (dreyer2023revealing; geirhos2018imagenet; geirhos2020shortcut; buolamwini2018gender; xiao2021noise; wang2019balanced; singh2020don). These include object-level attributes (e.g., color, material), contextual dependencies (e.g., background, co-occurring object state), and demographic associations. Together with an automated evaluation protocol, this benchmark provides a controlled environment for evaluating visual attribute reliance detection methods.
@@
Our method is model-agnostic and can be applied to any vision model that assigns scores to input images. To demonstrate its versatility, we evaluate it on both our controlled benchmark and state-of-the-art pretrained models. Across a range of visual reliance types, our experiments show that our self-reflective agent consistently outperforms non-reflective baselines. Moreover, it successfully uncovers previously unreported attribute dependencies in pretrained models (Fig. 1). For example, it identifies that the CLIP-ViT vision encoder (radford2021learning) recognizes teachers based on classroom backgrounds and that YOLOv8 (yolov8_ultralytics), trained for object detection for autonomous driving, relies on the presence of crosswalks to detect pedestrians. These findings highlight the efficacy of our method as a scalable tool for detecting hidden dependencies in pretrained models deployed in real-world scenarios.
@@ 2 Related Work
Revealing Visual Attribute Reliance in Vision Models.   Prior work has explored methods to uncover the visual cues that drive model predictions. One common strategy is to manipulate input features to isolate model sensitivities, such as shape or spectral biases in classifiers (gavrikov2024can), or attribute preferences in face recognition systems (liang2023benchmarking). Other works rely on interpretability tools to identify potential dependencies—for example, extracting keywords from captions of misclassified images (kim2024discovering), or using feature visualizations to reveal facial attribute reliance (teotia2022interpreting). However, most of these methods target specific types of biases and rely on predefined concept sets or human inspection. In contrast, our framework introduces a unified, flexible approach that can detect a broad range of attribute reliances without prior assumptions about the relevant features.
@@
Interpretability and Automated Analysis.   Initial work on interpretability automation produced textual descriptions of internal model features, using keywords (bau2017network), programs (mu2020compositional), or natural language summaries (hernandez2021natural; bills2023language; gandelsman2023interpreting). While informative, these descriptions are typically correlational, lacking behavioral validation (huang2023rigorously; schwettmann2023find; hausladen2024social). More recent work introduces agents that actively probe models. For instance, the Automated Interpretability Agent (AIA) (schwettmann2023find) used a language model to analyze black-box systems via a single pass over the input space. The Multimodal Automated Interpretability Agent (MAIA) (shaham2024multimodal) extended this approach by incorporating iterative experimentation and multimodal tools, enabling more detailed analysis of model internals. Our work builds on this direction by focusing specifically on discovering visual attribute reliances and introducing a self-reflection mechanism that allows the agent to revise faulty hypotheses based on experimental evidence, leading to more accurate and robust conclusions.
@@
Benchmarks for Visual Attribute Reliance Detection.   Standardized benchmarks for evaluating visual attribute reliance remain limited. Prior evaluations often use models trained on datasets with known biases, such as WaterBirds (wah2011caltech) or CelebA (liu2015deep), using label co-occurrence as a proxy for ground truth (sagawa2019distributionally). For generative settings, OpenBias (d2024openbias) proposes biases via LLMs, generates images from biased prompts, and assesses reliance using VQA models. However, OpenBias is not applicable to predictive models and cannot conduct controlled interventions. We introduce a suite of 130 vision models with explicitly injected attribute reliances across 18 categories, providing fine-grained control over attribute type and strength. This allows rigorous, scalable evaluation of reliance detection methods in a predictive setting.
@@
Agent-Based Reasoning and Self-Reflective Systems.   A growing body of work explores how agents can improve reasoning through reflection, feedback, and interaction. Methods like SELF-Refine (madaan2023selfrefine), Reflexion (shinn2023reflexion), and ReAct (yao2023react) introduce multi-step loops in which agents revise their outputs via self-critique. Similarly, Du et al. (du2023improving) show that multi-agent debate improves factual consistency and reasoning in language models. Our agent uses a task-specific self-evaluation protocol, enabling it to assess whether its conclusions align with actual model behavior. This integration of behavioral validation with self-reflection allows our agent to autonomously revise hypotheses and close the interpretability loop.
@@ 3.1 Hypothesis-Testing Stage
In this stage, SAIA iteratively refines hypotheses about the attribute sensitivities of the subject model. Inspired by MAIA (shaham2024multimodal), we design SAIA to operate in a scientific loop: it begins by proposing candidate attributes that the subject model might rely on, designs multiple experiments involving generating and editing images to test these hypotheses (e.g. edit an image with a suit to change its color), observes the resulting model’s behavior (e.g. measure the subject model scores across these experiments) and updates its beliefs accordingly. This cycle continues until SAIA converges on an initial conclusion about the model’s sensitivity to image features.
@@ 3.2 Self-Reflective Stage
Self-Evaluation protocol   Self-evaluation serves two key purposes: (i) to assess whether the current explanation matches the model’s behavior, and (ii) to guide further experimentation if it does not. The process begins by querying a separate language model instance (we use Claude-Sonnet-3.5) to generate two diverse sets of image prompts, that according to SAIA’s description are predicted to elicit high and low detection scores, respectively. The first set, termed the “predicted high-scoring images”, contains instances of images with the target concept (e.g., teddy bear) along with the detected reliance attribute (e.g., educational settings) that SAIA detected the model to be selective for. The second set, the “predicted low-scoring images”, contains instances of the target concept, but with the absence of the detected attribute. This prompt generation strategy emphasizes attribute-controlled diversity, where each prompt in the high or low-scoring group keeps the core attribute constant while allowing other visual factors to vary. These prompts are then used as inputs to a text-to-image model (we use Flux.1-dev flux2024) that generates the corresponding images, which are then fed to the subject model, and the output scores are recorded. If SAIA’s explanation is accurate, the model should exhibit systematically higher scores on the “predicted high-scoring” images and lower scores on the “predicted low-scoring” set.
Behavior-matching protocols of this kind have been shown to be effective in other evaluation settings, particularly for the task of producing textual labels of neurons’ behavior in pretrained models (shaham2024multimodal; kopf2024cosy; huang2023rigorously). In domains where ground-truth explanations are unavailable, they provide a way to validate hypotheses through behavioral consistency. We repurpose this evaluation method as a basis for self-reflection: SAIA uses it to validate its own conclusions, determine whether further experimentation is necessary, and reflect on its own findings based on measured model behavior.
@@ 4 A Benchmark of Models with Controlled Attribute Reliance
To evaluate the capabilities of SAIA, we constructed a benchmark of 130 unique object recognition models that exhibit 18 diverse types of visual attribute reliance. All simulated behaviors are inspired by known vulnerabilities of vision models (dreyer2023revealing; geirhos2018imagenet; geirhos2020shortcut; buolamwini2018gender; xiao2021noise; wang2019balanced; singh2020don), and mimic spurious correlations between the target object and image attributes such as object color, background context, co-occurring object state, or demographic cues. To assess the generalizability of our method, the benchmark also includes a subset of models with counterfactual attribute reliance that are intentionally rare or unnatural in real-world pretrained models (e.g., a suit detector responds more strongly when a women wear the suit). Each benchmark model includes an input parameter that controls the strength of the injected reliance, allowing for precise control over model behavior. Importantly, because these models are explicitly engineered with a known intended behavior, they serve as a controlled testbed for evaluating and comparing feature reliance detection methods.
@@ 4.1 Simulating attribute dependencies
Figure 4 illustrates a simulated attribute reliance scenario. Given a target object class $t$ (e.g. bird) and an intended injected attribute reliance $i$ (e.g. setting; beach), we simulate a model $\mathcal{C}_{t,i}$ that detects $t$ under the condition $i$. Each benchmark model is composed of two components; an object detector $\mathcal{O}_{t}$ and an attribute condition detector $\mathcal{A}_{i}$, which modulates the output of $\mathcal{O}_{t}$ based on the presence or absence of the specified attribute. To compute the final output of the model $\mathcal{C}_{t,i}$ on an input image img, we first pass the image through the object detector $\mathcal{O}_{t}$. If the target object class is not detected, the model returns a low random baseline score. If the object is detected, the image is then evaluated by the attribute condition detector $\mathcal{A}_{i}$. If the attribute condition is satisfied, the original subject model score $\mathcal{O}_{t}(\texttt{img})$ is returned, simulating full model response. Otherwise, the subject model score is discounted by a multiplication factor of $\alpha$, simulating attenuated confidence due to the missing attribute. The scalar $\alpha\in[0,1]$ controls the magnitude of the injected reliance: higher values of $\alpha$ simulate stronger reliance on the attribute. Please refer to Appendix B.2 for empirical evaluation of reliance magnitude as a function of $\alpha$. In all the benchmark models, we use Grounding DINO (liu2023grounding) as the object detector $\mathcal{O}_{t}$ and SigLIP (Zhai_2023_ICCV) as the attribute condition detector $\mathcal{A}_{i}$, which in practice is guided by a textual description of the injected attribute condition $i$. For demographic attribute dependencies, FairFace (karkkainenfairface) is used for $\mathcal{A}_{i}$ instead of SigLIP (see implementation details below). Notably, this model composition approach is highly flexible—one could engineer any object-condition pairing to construct an object detection model with a desired attribute reliance.
@@ 5.1 Evaluation protocol
Predictiveness score   Following kopf2024cosy; schwettmann2023find, we quantify how well a candidate reliance description matches model behavior. Similar to the self-evaluation score, given a candidate explanation, we start by generating 10 synthetic images that are expected to elicit high model scores and 10 that are expected to elicit low scores. We then pass these images through the model and record its actual responses. Each image is assigned a binary prediction label (high or low predicted response), and we threshold the model’s scores to obtain a binary outcome (high or low measured response). The predictiveness score is computed as the proportion of images where the predicted label matches the model’s actual binary output. This reflects how well the explanation predicts individual model responses.
LLM as a Judge   We use a language model as a judge (zheng2023judging) in a two-alternative forced choice (2AFC) setting. Given a ground-truth explanation of a benchmark model and two candidate explanations, one produced by SAIA and one from a baseline, the LLM judge is asked to choose which candidate better matches the ground-truth description. For each description pair, we repeat the test 10 times, and report the average preference rate for SAIA’s descriptions.
Baselines   We compare SAIA against the following alternatives: (i) Milan-style reliance detection: Following Milan (hernandez2021natural), this approach avoids iterative experimentation and detects the reliance based on a precomputed set of image exemplars that maximize the model’s scores.
@@
(ii) MAIA-style agent: Based on shaham2024multimodal, this method performs hypothesis testing but does not engage in self-reflection to revise its explanation. For a fair comparison, we equip MAIA with the same set of tools and backbone model that our method uses. (iii) Multi-agent ensemble: We run 10 independent MAIA-style, non-self-reflective agents and select the explanation with the highest predictiveness score. This tests whether repeated sampling alone can match the performance gains from self-reflection. (iv) Ground-truth descriptions: For benchmark models, we include ideal natural language descriptions of the injected reliance as an upper bound on performance.